# On the Design of One-step Diffusion via Shortcutting Flow Paths

### v3

@@ 4 Improvements to training
From the comparison between Fig. LABEL:fig:elucidscimgcnd and Fig. LABEL:fig:elucidscimgcfg, it is evident that classifier-free guidance (CFG) is crucial for high-quality image generation (meanflow). With CFG, the class-conditional velocity $\bm{v}_{t}(\bm{x}_{t}|\bm{x}_{0},c)$ leverages instance-level supervision from the label $c$. In contrast, $\bm{v}^{*}_{t}(\bm{x}_{t}|\{(\bm{y}^{(i)},c^{(i)})\}^{B}_{i=1}),$ is computed by averaging over randomly drawn mini-batches, which is likely to dilute or erase the class-specific signal. To this end, we employ a plug-in probability $p_{\text{plug-in}}$ that substitutes the conditional velocity with the plug-in velocity, as a trade-off between lowering variance during training and retaining class guidance. The other trick is class-consistent mini-batching: When applying CFG during training, we ensure that each mini-batch is sampled within the same class. In multi-GPU training, the class labels of mini-batches across different processes are independent of each other.
@@
Moreover, since sCT can be regarded as a variant of CTSC, several training strategies have already been explored in its original work, such as variational adaptive loss weighting (edm2) and tangent warmup (scm). These techniques are also applicable to CTSC and bring performance improvements in the cases given in Appendix D.3.
@@ 5 Scaling-up Evaluation
GAN BigGAN (biggan) 112M 1 6.95
GigaGAN (gigagan) 569M 1 3.45
StyleGAN-XL (stylegan) 166M 1 2.30
AR/Mask AR w/ VQGAN (ar) 227M 1024 26.52
MaskGIT (maskgit) 227M 8 6.18
@@
In this part, we evaluate the proposed ESC as an improved variant of CTSCs to illustrate its effectiveness at scale. We conduct a scaling-up experiment on ImageNet-$\times256$ in latent space, and employ SiT-XL/2 ($\sim$676M param.) as the backbone model. We follow the training setting of MeanFlow with CFG, where the model is trained from scratch with 240 epochs ($\sim$1.2M iterations). Furthermore, ESC+ is trained with 480 epochs ($\sim$2.4M iterations). In addition, for CIFAR-10 (cifar), all the shortcut models use the same U-Net (unet) architecture from sde ($\sim$55M param.). The code repository is provided for reproducibility [^1] https://github.com/EDAPINENUT/ExplicitShortCut/. For further details on setting, please refer to Appendix D.3.
@@ A.1 Stochastic Interpolants and Flow Map
Definition A.1 (Stochastic Interpolants (stochasticinterp)).
@@
Definition A.3 (Flow Map (flowmap; tutorial)).
@@
Proposition A.4 (Consistency Property (flowmap)).
@@ A.3 Derived Flow Path from preconditioner of EDM
the preconditioner of EDM is equivalent to the cosine path in our paper, or namely TrigFlow in sCT, by using the change-of-variable. This part is mostly based on Appendix B of TrigFlow proposed by scm, while we use a more unified view from stochastic interpolants (stochasticinterp) and SiT (sit).
@@
We draw $\bm{x}_{0}\sim p_{\text{data}}$ and $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ and define
@@
with scaling coefficients ensuring unit-normalized training targets:
@@
A.3.2 From EDM preconditioner to cosine path.
To further simplify, introduce a normalized representation
$\gamma_{t}=\frac{\alpha_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}},\qquad\beta_{t}=\frac{\sigma_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}},\qquad\hat{\bm{x}}_{t}=\frac{\bm{x}_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}=\gamma_{t}\bm{x}_{0}+\beta_{t}\bm{\varepsilon},$ (36)
so that $\gamma_{t}^{2}+\beta_{t}^{2}=1$. In this normalized form, the coefficients become
$c_{\mathrm{in}}(t)=\frac{1}{\sigma_{\text{data}}},\qquad c_{\mathrm{skip}}(t)=\gamma_{t},\qquad c_{\mathrm{out}}(t)=-\,\sigma_{\text{data}}\,\beta_{t},$ (37)
@@
${D}^{\theta}(\bm{x}_{t},t)=\gamma_{t}\,\hat{\bm{x}}_{t}-\beta_{t}\,\sigma_{\text{data}}\,\varphi^{\theta}\!\big(\hat{\bm{x}}_{t}/\sigma_{\text{data}},\,c_{\mathrm{noise}}(t)\big).$ (38)
@@
Since $\gamma_{t}^{2}+\beta_{t}^{2}=1$, we can introduce a cosine time variable $t^{\prime}\in[0,1]$ such that
$\gamma_{t}=\cos\!\Big(\tfrac{\pi}{2}t^{\prime}\Big),\qquad\beta_{t}=\sin\!\Big(\tfrac{\pi}{2}t^{\prime}\Big),\qquad t^{\prime}=\tfrac{2}{\pi}\arctan\!\Big(\tfrac{\beta_{t}}{\gamma_{t}}\Big)=\tfrac{2}{\pi}\arctan\!\Big(\tfrac{\sigma(t)}{\sigma_{\text{data}}}\Big).$ (39)
@@ A.3.3 From Score parameterization to Velocity
Since ${D}^{\theta}(\bm{x}_{t},t)$ aims to approximate
$F^{\theta}(\bm{x}_{t},t)=\frac{\pi}{2}\,\sigma_{\text{data}}\,{\varphi}^{\theta}\!\big(\bm{x}_{t}/\sigma_{\text{data}},\,c^{\prime}_{\text{noise}}(t)\big),\qquad v_{t}^{\theta}(\bm{x}_{t})=F^{\theta}(\bm{x}_{t},t).$
the coefficient $\cos(\frac{\pi}{2}t)$ and $\frac{2}{\pi}\sin(-\frac{\pi}{2}t)$ coincides to $\bar{\alpha}(t,0)$ and $\bar{\beta}(t,0)$ in Eq. 28, the parameterization of $v_{t}^{\theta}(\bm{x}_{t})=F^{\theta}(\bm{x}_{t},t)$ is equivalent to denoise the path of preconditioner in EDM schedule. The same evidence is also provided with Eq. 4 in scm.
@@ B.1 Consistency Training
In CTs and CMs (cm), the original paper uses the EDM preconditioner as the components of the basic diffusion model. By using a neural network $\varphi^{\theta}$ to approximate the score function $\bm{s}_{t}(\bm{x})$, its target is to map any noised samples in $t$ to 0 which in the flow map notation, reads
@@ B.2 Shortcut Diffusion
In the original paper of SCD (scd), it parameterizes the velocity with the neural network as
@@ B.3 Inductive Moment Matching
Further, we bridge the IMM loss with the common flow map learning objective in the following. In Eq. 52, it gives the group kernel function, according to Appendix. C.3 in imm, we can write the loss here as
@@ B.4 MeanFlow
However, the marginal velocity $\bm{v}_{t}$ is inaccessible in training, so it can be replaced by the conditional velocity $\bm{v}_{t|0}$. From Eq. 6 in meanflow, by adding the adaptive loss term $w$, this loss coincides with the training objective of MeanFlow in Eq. 8, as
@@ Appendix C Proof of Theorems and Propositions
We note that MeanFlow loss and sCT loss are all $\mathcal{L}_{\text{ctsc}}(\theta)$, and CT loss and SCD loss are all $\mathcal{L}_{\text{dtsc}}(\theta)$. IMM’s loss is calculated across different conditional paths, as finally bounded by the $\mathcal{L}_{\text{dtsc}}(\theta)$ as shown in Appendix B.3. The mentioned previous methods all satisfy the assumptions about $w(r,s,t)$, $p(r,s,t)$, and $\hat{X}_{t,t}(\bm{x}_{t})=\bm{x}_{t}$. As for the assumption of $d(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|_{2}^{2}$, it holds for all the mentioned methods except CT, which takes LPIPS as the metric function. The convergence of $\mathcal{L}_{\text{ct}}$ has already been proved by cm.
@@ C.2.2 Error Bound of Start Point Differential CTSC
Next, we prove the error bound of the start point differential CTSC. The derivation is adopted from flowmap.
@@ C.2.3 Error Bound of End Point Differential CTSC
Previous works have claimed that the cosine path is optimal for diffusion models from the perspective of Fisher information metric (santos2023usingornsteinuhlenbeckprocessunderstand). Here, we provide the analysis of the optimal path for one-step models under the Fisher information metric.
@@
the Fisher information metric becomes (Karczewski2025SpacetimeGO)
@@
(zhang2025cosine) The optimal schedule under the metric $I(\gamma)\in\mathbb{R}$ is generated by $\varphi^{*}$ of the form
@@ C.5.2 The Bias and Variance of Plug-in Velocity
The previous work (Kloeckner2018EmpiricalMR) has proven that
@@
We point out that the Wasserstein-1 distance between the empirical distribution and the true distribution decreases as the number of data samples increases, which has been established in previous literature (fournier2015rate).
@@ D.2 Details for Empirical Analysis of Table 2
sCM training techniques. In addition, in ESC, we, following sCM (scm) and EDMv2 (edm2), introduce a variational weighting output head, where the output of the time embedder is passed through a linear layer to a one-dimensional scalar as the adaptive weighting function, which reads $w_{\text{adpt}}^{\psi}(t,r)$ and is then used to reweight the original loss in Eq. 8. We keep the the SiT architecture blocks untouched, while architectural improvements are orthogonal and possible. Moreover, a ratio $r_{\text{grad}}=\min\{\frac{\text{iter}}{K_{\text{grad}}},1\}$ for tangent warmup is implemented for mitigating some gradient spikes during training, where $K_{\text{grad}}$ is set as 10k, the same as sCT training.
@@
optimizer Adam (kingma2017adammethodstochasticoptimization) Adam
@@
We conduct class-unconditional generation experiments on CIFAR-10. Following the official MeanFlow setting, we adopt the Adam optimizer with a learning rate of $6\times 10^{-4}$, batch size $1024$, and momentum parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$. We use a dropout rate of $0.2$, no weight decay, and an EMA decay factor of $0.99995$. Training is performed for 800k iterations, including a 10k warm-up phase. For time sampling, we draw $(r,t)\sim\mathrm{LogNorm}(-2.0,\,2.0)$, with probability $75\%$ that $r\neq t$. The adaptive weighting exponent is set to $p=0.75$. Data augmentation follows the protocol of edm, except that vertical flipping and rotation are disabled.
@@ D.4 Visualization Examples for ESC
GAN BigGAN (biggan) 112M 1 6.95
GigaGAN (gigagan) 569M 1 3.45
StyleGAN-XL (stylegan) 166M 1 2.30
AR/Mask AR w/ VQGAN (ar) 227M 1024 26.52
MaskGIT (maskgit) 227M 8 6.18
@@
Distill Diff-Instruct (diffinstruct) 1 4.53
DMD (dmd) 1 2.66
SID (sid) 1 1.92
Shortcut iCT (icm) 1 2.83
ECT (geng2025consistency) 1 3.60
sCT (scm) 1 2.97
IMM (imm) 1 3.20
MeanFlow (meanflow) 1 2.92
@@ E.2 Large Models Gain More Performance from Low Variance Training
  • Statistical Generalization. As the parameter space dimensionality increases, gradient noise also grows, making variance-reduction mechanisms—such as EMA, momentum, gradient clipping, or the proposed plug-in velocity—more beneficial. This observation aligns with the theoretical intuition in kaplan2020scalinglawsneurallanguage, where the generalization gap (or overfitting) is linked to the variance term scaling. Within the scaling law framework, bias dominates in smaller models, while variance becomes the main factor as the model scales up. To illustrate this, we compare the FID convergence curves of ESC-B/2 vs. MeanFlow-B/2 (trained for 600k iterations) and ESC-XL/2 vs. MeanFlow-XL/2 (trained for 1.2M iterations), as shown in Figure 11(c). Empirically, in the smaller B/2 setting, both methods converge rapidly to similar FID values. However, in the larger XL/2 model, MeanFlow’s FID curve plateaus in the later training stages, while ESC continues to improve and reaches 2.85. This suggests that in large-scale models, variance dominates generalization behavior, and the variance reduction introduced by plug-in velocity significantly enhances final performance.

### v5

@@ 4 Improvements to training
From the comparison between Fig. LABEL:fig:elucidscimgcnd and Fig. LABEL:fig:elucidscimgcfg, it is evident that classifier-free guidance (CFG) is crucial for high-quality image generation (Geng et al., 2025a). With CFG, the class-conditional velocity $\bm{v}_{t}(\bm{x}_{t}|\bm{x}_{0},c)$ leverages instance-level supervision from the label $c$. In contrast, $\bm{v}^{*}_{t}(\bm{x}_{t}|\{(\bm{y}^{(i)},c^{(i)})\}^{B}_{i=1}),$ is computed by averaging over randomly drawn mini-batches, which is likely to dilute or erase the class-specific signal. To this end, we employ a plug-in probability $p_{\text{plug-in}}$ that substitutes the conditional velocity with the plug-in velocity, as a trade-off between lowering variance during training and retaining class guidance. The other trick is class-consistent mini-batching: When applying CFG during training, we ensure that each mini-batch is sampled within the same class. In multi-GPU training, the class labels of mini-batches across different processes are independent of each other.
@@
Moreover, since sCT can be regarded as a variant of CTSC, several training strategies have already been explored in its original work, such as variational adaptive loss weighting (Karras et al., 2024) and tangent warmup (Lu and Song, 2025). These techniques are also applicable to CTSC and bring performance improvements in the cases given in Appendix D.3.
@@ 5 Scaling-up Evaluation
GAN BigGAN (Brock et al., 2019) 112M 1 6.95
GigaGAN (Kang et al., 2023) 569M 1 3.45
StyleGAN-XL (Karras et al., 2019) 166M 1 2.30
AR/Mask AR w/ VQGAN (Esser et al., 2021) 227M 1024 26.52
MaskGIT (Chang et al., 2022) 227M 8 6.18
@@
In this part, we evaluate the proposed ESC as an improved variant of CTSCs to illustrate its effectiveness at scale. We conduct a scaling-up experiment on ImageNet-$\times256$ in latent space, and employ SiT-XL/2 ($\sim$676M param.) as the backbone model. We follow the training setting of MeanFlow with CFG, where the model is trained from scratch with 240 epochs ($\sim$1.2M iterations). Furthermore, ESC+ is trained with 480 epochs ($\sim$2.4M iterations). In addition, for CIFAR-10 (Krizhevsky, 2009), all the shortcut models use the same U-Net (Ronneberger et al., 2015) architecture from Song et al. (2020) ($\sim$55M param.). The code repository is provided for reproducibility [^1] https://github.com/EDAPINENUT/ExplicitShortCut/. For further details on setting, please refer to Appendix D.3.
@@ A.1 Stochastic Interpolants and Flow Map
Definition A.1 (Stochastic Interpolants (Albergo et al., 2023)).
@@
Definition A.3 (Flow Map (Boffi et al., 2025a; Liu, 2025)).
@@
Proposition A.4 (Consistency Property (Boffi et al., 2025a)).
@@ A.3 Derived Flow Path from preconditioner of EDM
the preconditioner of EDM is equivalent to the cosine path in our paper, or namely TrigFlow in sCT, by using the change-of-variable. This part is mostly based on Appendix B of TrigFlow proposed by Lu and Song (2025), while we use a more unified view from stochastic interpolants (Albergo et al., 2023) and SiT (Ma et al., 2024).
@@
We draw $\bm{x}_{0}\sim p_{\text{data}}$ and $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\sigma_{\text{data}}\bm{I})$ and define
@@
since $\frac{\sigma_{\text{data}}^{2}}{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}(\bm{x}_{0}+\sigma(t)\bm{\varepsilon})=c_{\text{skip}}(t)\bm{x}_{t}$, and according to Eq. 29 and Eq. 30, with scaling coefficients ensuring unit-normalized training targets:
@@
${D}^{\theta}(\bm{x}_{t},t)=\alpha_{t}\,\hat{\bm{x}}_{t}-\beta_{t}\,\sigma_{\text{data}}\,\varphi^{\theta}\!\big(\hat{\bm{x}}_{t}/\sigma_{\text{data}},\,c_{\mathrm{noise}}(t)\big).$ (36)
@@
Since $\alpha_{t}^{2}+\beta_{t}^{2}=1$, we can introduce a cosine time variable $t^{\prime}\in[0,1]$ such that
$\alpha_{t}=\cos\!\Big(\tfrac{\pi}{2}t^{\prime}\Big),\qquad\beta_{t}=\sin\!\Big(\tfrac{\pi}{2}t^{\prime}\Big),\qquad t^{\prime}=\tfrac{2}{\pi}\arctan\!\Big(\tfrac{\beta_{t}}{\alpha_{t}}\Big)=\tfrac{2}{\pi}\arctan\!\Big(\tfrac{\sigma(t)}{\sigma_{\text{data}}}\Big).$ (37)
@@ A.3.3 From Score parameterization to Velocity
Since ${D}^{\theta}(\bm{x}_{t},t)$ aims to approximate $\bm{x}_{0}$, and if we write
$\bm{v}^{\theta}(\bm{x}_{t},t)=\frac{\pi}{2}\,\sigma_{\text{data}}\,{\varphi}^{\theta}\!\big(\bm{x}_{t}/\sigma_{\text{data}},\,c^{\prime}_{\text{noise}}(t)\big).$
, it reads
${D}^{\theta}(\bm{x}_{t},t)=\cos(\frac{\pi}{2}t){\bm{x}}_{t}+\frac{2}{\pi}\sin(-\frac{\pi}{2}t)\bm{v}^{\theta}(\bm{x}_{t},t)$
the coefficient $\cos(\frac{\pi}{2}t)$ and $\frac{2}{\pi}\sin(-\frac{\pi}{2}t)$ coincides to $\bar{\alpha}(t,0)$ and $\bar{\beta}(t,0)$ in Eq. 28, the parameterization of $\bm{v}_{t}^{\theta}(\bm{x}_{t})=\frac{\pi}{2}\sigma_{\text{data}}F^{\theta}(\bm{x}_{t},t)$ is equivalent to denoise the path of preconditioner in EDM schedule. The same evidence is also provided with Eq. 4 in Lu and Song (2025).
@@ B.1 Consistency Training
In CTs and CMs (Song et al., 2023), the original paper uses the EDM preconditioner as the components of the basic diffusion model. By using a neural network $\varphi^{\theta}$ to approximate the score function $\bm{s}_{t}(\bm{x})$, its target is to map any noised samples in $t$ to 0 which in the flow map notation, reads
@@ B.2 Shortcut Diffusion
In the original paper of SCD (Frans et al., 2025), it parameterizes the velocity with the neural network as
@@ B.3 Inductive Moment Matching
Further, we bridge the IMM loss with the common flow map learning objective in the following. In Eq. 50, it gives the group kernel function, according to Appendix. C.3 in Zhou et al. (2025), we can write the loss here as
@@ B.4 MeanFlow
However, the marginal velocity $\bm{v}_{t}$ is inaccessible in training, so it can be replaced by the conditional velocity $\bm{v}_{t|0}$. From Eq. 6 in Geng et al. (2025a), by adding the adaptive loss term $w$, this loss coincides with the training objective of MeanFlow in Eq. 8, as
@@ Appendix C Proof of Theorems and Propositions
We note that MeanFlow loss and sCT loss are all $\mathcal{L}_{\text{ctsc}}(\theta)$, and CT loss and SCD loss are all $\mathcal{L}_{\text{dtsc}}(\theta)$. IMM’s loss is calculated across different conditional paths, as finally bounded by the $\mathcal{L}_{\text{dtsc}}(\theta)$ as shown in Appendix B.3. The mentioned previous methods all satisfy the assumptions about $w(r,s,t)$, $p(r,s,t)$, and $\hat{X}_{t,t}(\bm{x}_{t})=\bm{x}_{t}$. As for the assumption of $d(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|_{2}^{2}$, it holds for all the mentioned methods except CT, which takes LPIPS as the metric function. The convergence of $\mathcal{L}_{\text{ct}}$ has already been proved by Song et al. (2023).
@@ C.2.2 Error Bound of Start Point Differential CTSC
Next, we prove the error bound of the start point differential CTSC. The derivation is adopted from Boffi et al. (2025a).
@@ C.2.3 Error Bound of End Point Differential CTSC
Previous works have claimed that the cosine path is optimal for diffusion models from the perspective of Fisher information metric (Santos and Lin, 2023). Here, we provide the analysis of the optimal path for one-step models under the Fisher information metric.
@@
the Fisher information metric becomes (Karczewski et al., 2025)
@@
(Zhang, 2025) The optimal schedule under the metric $I(\gamma)\in\mathbb{R}$ is generated by $\varphi^{*}$ of the form
@@ C.5.2 The Bias and Variance of Plug-in Velocity
The previous work (Kloeckner, 2018) has proven that
@@
We point out that the Wasserstein-1 distance between the empirical distribution and the true distribution decreases as the number of data samples increases, which has been established in previous literature (Fournier and Guillin, 2015).
@@ D.2 Details for Empirical Analysis of Table 2
sCM training techniques. In addition, in ESC, we, following sCM (Lu and Song, 2025) and EDMv2 (Karras et al., 2024), introduce a variational weighting output head, where the output of the time embedder is passed through a linear layer to a one-dimensional scalar as the adaptive weighting function, which reads $w_{\text{adpt}}^{\psi}(t,r)$ and is then used to reweight the original loss in Eq. 8. We keep the the SiT architecture blocks untouched, while architectural improvements are orthogonal and possible. Moreover, a ratio $r_{\text{grad}}=\min\{\frac{\text{iter}}{K_{\text{grad}}},1\}$ for tangent warmup is implemented for mitigating some gradient spikes during training, where $K_{\text{grad}}$ is set as 10k, the same as sCT training.
@@
optimizer Adam (Kingma and Ba, 2017) Adam
@@
We conduct class-unconditional generation experiments on CIFAR-10. Following the official MeanFlow setting, we adopt the Adam optimizer with a learning rate of $6\times 10^{-4}$, batch size $1024$, and momentum parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$. We use a dropout rate of $0.2$, no weight decay, and an EMA decay factor of $0.99995$. Training is performed for 800k iterations, including a 10k warm-up phase. For time sampling, we draw $(r,t)\sim\mathrm{LogNorm}(-2.0,\,2.0)$, with probability $75\%$ that $r\neq t$. The adaptive weighting exponent is set to $p=0.75$. Data augmentation follows the protocol of Karras et al. (2022), except that vertical flipping and rotation are disabled.
@@ D.4 Visualization Examples for ESC
GAN BigGAN (Brock et al., 2019) 112M 1 6.95
GigaGAN (Kang et al., 2023) 569M 1 3.45
StyleGAN-XL (Karras et al., 2019) 166M 1 2.30
AR/Mask AR w/ VQGAN (Esser et al., 2021) 227M 1024 26.52
MaskGIT (Chang et al., 2022) 227M 8 6.18
@@
Distill Diff-Instruct (Luo et al., 2024) 1 4.53
DMD (Yin et al., 2024b) 1 2.66
SID (Zhou et al., 2024) 1 1.92
Shortcut iCT (Song and Dhariwal, 2023) 1 2.83
ECT (Geng et al., 2025b) 1 3.60
sCT (Lu and Song, 2025) 1 2.97
IMM (Zhou et al., 2025) 1 3.20
MeanFlow (Geng et al., 2025a) 1 2.92
@@ E.2 Large Models Gain More Performance from Low Variance Training
  • Statistical Generalization. As the parameter space dimensionality increases, gradient noise also grows, making variance-reduction mechanisms—such as EMA, momentum, gradient clipping, or the proposed plug-in velocity—more beneficial. This observation aligns with the theoretical intuition in Kaplan et al. (2020), where the generalization gap (or overfitting) is linked to the variance term scaling. Within the scaling law framework, bias dominates in smaller models, while variance becomes the main factor as the model scales up. To illustrate this, we compare the FID convergence curves of ESC-B/2 vs. MeanFlow-B/2 (trained for 600k iterations) and ESC-XL/2 vs. MeanFlow-XL/2 (trained for 1.2M iterations), as shown in Figure 11(c). Empirically, in the smaller B/2 setting, both methods converge rapidly to similar FID values. However, in the larger XL/2 model, MeanFlow’s FID curve plateaus in the later training stages, while ESC continues to improve and reaches 2.85. This suggests that in large-scale models, variance dominates generalization behavior, and the variance reduction introduced by plug-in velocity significantly enhances final performance.